This is the last portion of my homework and my tutor is out of town for the holidays, is it possible that someone can assist me on this? Thank you!
Determine whether the following equations have a solution or not? Justify your answer.
a) x2 + 6x - 7 = 0
b) z2 + z + 1 = 0
c) (3)1/2y2 - 4y - 7(3)1/2 = 0
d) 2x2 - 10x + 25 = 0
e) 2x2 - 6x + 5 = 0
f) s2 - 4s + 4 = 0
g) 5/6x2 - 7x - 6/5 = 0
h) 7a2 + 8a + 2 = 0

Answer by checkley77(12844)

a) x2 + 6x - 7 = 0
(x+7)(x-1)=0
x+7=0
x=-7 ans.
x-1=0
x=1 ans.
-------------------------
b) z2 + z + 1 = 0
z = (-b +- sqrt( b^2-4*a*c ))/(2*a) }}
z=(-1+-sqrt[1^2-4*1*1])/2*1
z=(-1+-sqrt[1-4])/2
z=(-1+-sqrt-3)/2
z=(-1+-1.732)/2
z=-1/2+-.866i
----------------------------------------
I'll leave the rest for you to solve.
c) (3)1/2y2 - 4y - 7(3)1/2 = 0
d) 2x2 - 10x + 25 = 0
e) 2x2 - 6x + 5 = 0
f) s2 - 4s + 4 = 0
g) 5/6x2 - 7x - 6/5 = 0
h) 7a2 + 8a + 2 = 0